OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −1

OpenStudy (anonymous):

Set the first equation = to. y = -2x+ 1 Now substitute it for y into equation 2 and solve for x.

OpenStudy (anonymous):

Get the first equation equal to y so y=1-2x. Then substitute that into the second equation making y=1-2x. That should give you 4x + 2(1-2x) = -1

OpenStudy (anonymous):

x can be any value as there is an infinite number of solutions.

OpenStudy (anonymous):

@ivandelgado, no it cannot.

OpenStudy (anonymous):

There is no x value as there is no solution.

OpenStudy (anonymous):

There is a solution. Did you even read what we told you?

OpenStudy (anonymous):

If there are 2 equations, where x and y are the same in each, there has to be a solution.

OpenStudy (anonymous):

If you can find me two solutions for x, I will pay you $1000.

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

@alrightatmaths, what are you talking about??? X has 1 solution. Y has one solution in the porblem here

OpenStudy (anonymous):

@Zelda @alrightatmaths

OpenStudy (anonymous):

I know, that's what I just said.

OpenStudy (anonymous):

I said there has to be a solution. And I was saying that he couldn't find 2 values for x.

OpenStudy (anonymous):

so its 2

OpenStudy (anonymous):

There are no solutions to that.

OpenStudy (anonymous):

@ivandelgado Oh, wow, I'm sorry, your first guess was right. There are no solutions as the lines are parallel.

OpenStudy (anonymous):

Yeah, just realized the same thing :P